Please help with my home work
1. solve by factoring
x^2-10x = -16
2. Solve using the square root property
5x^2 -1 =74
3. Solve using the square root property
(x + 6)^2 = 121

Answer by checkley77(12844)

x^2-10x = -16
x^2-10x+16=0
(x-8)(x-2)=0
x-8=0
x=8 answer.
x-2=0
x=2 answer.
------------------------------
5x^2 -1 =74
5x^2-1-74=0
5x^2-75=0
5(x^2-25)=0
5(x-5)(x+5)=0
x-5=0
x=5 answer.
x+5=0
x=-5 answer.
-----------------------------
(x + 6)^2 = 121
x^2+12x+36-121=0
x^2+12x-85=0
(x+17)(x-5)=0
x+17=0
x=-17 answer.
x-5=0
x=5 answer.